OpenStudy (anonymous):

Find the number a such that the line x = a divides the region bounded by the curves y = x, y = 0, and x = 4 into two regions with equal area. Give your answer correct to 3 decimal places.

OpenStudy (danjs):



OpenStudy (danjs):

set up a couple of equations with similar triangles and areas, or use an integral 0 to a = a to 4, maybe

OpenStudy (dumbcow):

note that because of line "y=x" ---> h = a Area of small triangle = 1/2 Area of big triangle \[\frac{1}{2} a^2 = \frac{1}{2}(\frac{1}{2} 4^2)\] \[\frac{1}{2} a^2 = 4\] \[a = \sqrt{8} = 2 \sqrt{2}\]